# Identifying the Truth of Global Model: A Generic Solution to Defend Against Byzantine and Backdoor Attacks in Federated Learning (full version)

### v2

@@ 3.5 Proof of Byzantine-Resilience
Because $S$ is an arbitrary set of size $n-f$, we say that
@@ 3.6 Resisting against Adaptive Attack on FedTruth
The effectiveness of FedTruth relies on the assumption that the majority of clients are reliable and diverse. If an adversary compromises more than 50% of the clients, they can dominate the results of FedTruth. From our experimental results, we find that when an adversary compromises 40% (4 out of 10) clients in each round, FedTruth can still prevent Byzantine attacks, as seen in Figure 2. However, when the non-iid degree is further increased to 95%, as shown in Figure 7, the accuracy drops and convergence speed becomes slow because some uncompromised clients may perform poorly with highly non-iid training data, leading to a bad estimation of the ground truth by FedTruth. However, our results outperformed all other aggregation algorithms during this experiment, excluding FLTrust.
@@ 3.7 Deploying FedTruth in Each Layer?
One major challenge in truth discovery is data heterogeneity, which may include non-structured data and missing values. However, this challenge is not applicable to FedTruth because all the local model updates are in the same structure. For example, in deep neural networks, the model updates can be represented as multiple-layer tensors. FedTruth can be run for just one time by the aggregator to compute the truth of model updates by feeding all the local model updates into the FedTruth algorithm. This deployment treats the local model update as an observed value in the truth discovery algorithms. Also, we can deploy FedTruth on each layer to estimate the truth of that single layer, which means that the weights allocated to all the clients may vary on different layers. We denote this deployment as FedTruth-layer. Such layer-wise deployment seems reasonable, it also brings the computation overhead which is linear to the number of layers. We compare the efficiency between FedTruth and FedTruth-layer in the Section 4.6.
@@ 4.1 Experimental Settings
Datasets: We conduct the experiments with MNIST [7], FMNIST [30], and CIFAR-10 [16] datasets. FMNIST and MNIST are comprised of 60,000 black-and-white labeled images of size (28??28). MNIST contains handwritten digits, and FMNIST consists of images of clothing items. CIFAR-10 consists of 60000 color images of size (32x32). During the Edge-Case attack experiment, we used the Arkiv Digital Sweden (ARDIS) [17] dataset as the adversarial backdoor images. The ARDIS dataset consists of handwritten digits originating from Swedish church records. This dataset is suitable for targeted images when inserting backdoors into MNIST, as ARDIS entirely consists of naturally occurring edge cases.
Clients: When crafting the clients’ local datasets, we draw their datapoints randomly from a non-iid distribution. We use a non-iid distribution because it better represents clients’ data in practice than an iid distribution. The client’s local data is generated a non-iid distribution, where the bias parameter default is $0.8$. In addition, we evaluate the impact of non-iid degree using the model-boosting attacks, as seen in Section 4.4. In each FL round, we randomly select 10 clients and choose a subset of these selected clients as adversarial clients.
@@ 4.2 Byzantine Attacks
Figure 2(a) shows how all of the aggregation algorithms perform when there are no adversaries present. Figure 2(b) presents the results when $10\%$ of the clients are compromised in each round. We observe that all of the aggregation algorithms performed well except for FedAvg. However, when the percentage of compromised clients increases to 30%, as shown in Figure 2(c), the FedTruth methods remain unaffected by the attack, regardless of the number of adversaries. However, Trimmed-mean, similar to FedAvg, is significantly impacted at this stage and does not reach convergence.
The results in Figure 2(d) show that increasing the number of adversaries per iteration to 40% slows the convergence rate for the FedTruth, FedTruth-layer, and Median aggregation algorithms. However, they are still able to reach an accuracy of 80% after the 100th iteration. The FedAvg and Trimmed-mean algorithms were compromised during this version of the experiment as well, preventing them both from reaching any convergence when at least $20\%$ of the clients are adversarial. In contrast, the remaining algorithms (FLTrust, Krum, FLAME) were not affected during this attack regardless of the number of adversaries we selected.
Gaussian Noise Attack: The Gaussian noise attack aims to degrade the performance of the global model by adding Gaussian noise to the model. The noise is drawn from a multivariate Gaussian distribution $N(0,\sigma^{2}I)$ [5, 4] and is added directly to the adversaries’ model before sending it to the aggregator.
@@
In Figure 3(a), we show the accuracy and convergence speed of the model for all the aggregation algorithms against the Gaussian noise attack, where three adversaries launch this attack per round. Our findings are as follows: 1) FedTruth can defend against the Gaussian noise attack without significantly slowing down the convergence speed; 2) FedTruth can achieve the same model accuracy as FLTrust, which requires a benign dataset, showing that our FedTruth can actually find the ground truth of the model updates; 3) The convergence speed is impacted in FLAME and Krum, and the model accuracy of FLAME and Krum is not as high as that of FedTruth and FLTrust; and 4) FedAvg cannot converge within 100 rounds under the Gaussian noise attack.
@@ 4.3 Backdoor Attacks
In this section, we present our finding for target task accuracy (accuracy on an adversarial backdoor dataset) and the main task accuracy (accuracy on a benign dataset) for the distributed backdoor, projected gradient descent, and edge case attacks observed during each round of aggregation. We also present our findings on the effect that different distance functions have on these attacks.
@@
Figures 4(a), 5(a), and 6(a) show the main task accuracy for all of the versions of the DBA attack. Here, we see that both Krum and Flame are unable to train the main task during the base attack and when combined with the model-boosting. We suspect these results could be caused by the large number of adversaries colluding during this experiment or the imbalanced sampled data. The remaining algorithms are able to reach convergence on the main task during these attacks. When the DBA is combined with the constrain-and-scale attack, all algorithms are able to reach convergence, including Flame and Krum, which we suspect is due to the constrain-and-scale attack scaling the amount that an adversarial model can diverge during each epoch. However, FedTruth and FedTruth-layer experience a slower convergence rate during the DBA with model-boosting and constrain-and-scale attacks.
@@
Figures 4(b), 5(b), and 6(b) present our findings on the targeted task accuracy (i.e., the backdoor accuracy) during the DBA experiments. The adversarial goal during this attack is to add the targeted artifact to the global model without being detected and not affecting the model’s performance on the main task. Therefore, we will not be considering the Krum and FLAME algorithms during the base and model-boosting attacks, as they were unable to converge when training the main task, as seen in Figure 4(a). FedTruth in Figure 4(b) is the only aggregation algorithm that is not excluded that exceeds the 60% accuracy after the 60th FL iteration. We suspect that this is due to an inverse relationship with the effects of flattening the layers when subject to stealthy attacks as the number of layers in the model increases. Furthermore, during the DBA with model-boosting attack, we observe similar results for the algorithms that we are considering, with the exception that FedAvg exceeds our threshold while FedTruth can remove the targeted task. This is a result of a tradeoff between the stealthiness of the base DBA attack being diminished when increasing the amplitude of the adversarial updates vector in hopes of replacing the global model with the adversarial model. Our results also show that the DBA with constrain-and-scale attack is effective at adding the targeted task into FedAvg, FedTruth, Median, and Trimmed mean, reaching a backdoor accuracy above the 40% threshold after 100 iterations. However, FedTruth-layer, Krum, and FLTrust are able to remove the adversarial artifact.
@@ 4.4 The Impact of non-iid on FedTruth
Figure 7 shows how various non-iid bias parameters affect the experiments when adversaries apply the model-boosting attack. During these experiments, three adversaries were selected in each round. Our methodology for sampling non-iid data was specifically engineered to replicate varying degrees of label bias, thus enabling an in-depth analysis of its influence on federated learning model efficacy. We manipulated a bias parameter to adjust the label proportions within each client’s local dataset. For instance, setting the bias parameter to 0.9 indicated that 90% of a client’s dataset contained instances of their primary label, with the remaining 10% consisting of instances from other labels, allocated based on a Gaussian distribution. For this experiment, we set the bias parameters as $0.1$, $0.3$, $0.5$, $0.8$, and $0.95$.
@@ 4.5 Distance Function in FedTruth
Figures 8(a) and  8(b) show how the choice of distance metrics affects FedTruth and FedTruth-layer during Byzantine attacks. As expected, the two metrics that determine the difference between two points (Euclidean and Manhattan distances) performed the best, reaching convergence in both cases. This is because these approaches can easily determine the adversarial model furthest from the benign models as they are boosted or have additional random noise inserted into them. However, since the angular distance and cosine distance metrics do not consider a model’s magnitude, it is ineffective during the model boosting attack Figures 8(a). Additionally, since the Gaussian noise attack will only slightly modify the angle, it is not completely effective at removing the adversarial models when using angular distance and cosine distance.
Figures 8(c) and  8(d) present our results for all our distance metrics during a backdoor (DBA) attack. These results show that the inverse-cosine similarity or angular distance metrics perform the best as they are able to remove any backdoors injected into the model. However, our main and backdoor accuracy results look good for the models that determine the distance between two points. The custom distance unexpectedly cannot properly train the main task during this experiment.
More results on different distance metrics of FedTruth will be shown in the appendix.
@@ 4.6 Efficiency Evaluation of FedTruth and FedTruth-layer
We find that during the MNIST experiments, the Median algorithm is the most efficient, while Trimmed mean, FedAvg, and FLTrust take less than 0.01 seconds. FLAME takes about 0.16 seconds to filter and clip, while Krum takes around 0.08 seconds. Interestingly, FedTruth has a slightly faster aggregation time than FLAME, and FedTruth-layer is the slowest aggregation algorithm with an average aggregation time of 0.7 seconds.
@@
We count the number of iterations required for both algorithms to reach convergence and present the results in Table 2. For the CNN (8 layers) model on the MNIST dataset, we find that FedTruth requires an average of 5.17 iterations to estimate the ground-truth model update, while FedTruth-layer requires an average of 33.73 iterations (six times more than FedTruth) to reach convergence, despite having eight layers in the CNN model. Moreover, FedTruth on each layer has a smaller input size and requires less computation time on the distance compared to FedTruth with the entire model update as input. In conclusion, we find that both FedTruth and FedTruth-layer have similar performance on the MNIST CNN (8 layers) model regarding accuracy and robustness. However, FedTruth is much more efficient than FedTruth-layer when there are a large number of layers in the CIFAR-10 (ResNet-18) model.
We further evaluate the average aggregation time for 10, 100 and 1000 clients in a single FL round in Table 3, where the aggregation time is calculated as the average of 100 FL rounds. We can see that FedTruth and FedTruth-layer are as efficient as FLTrust (which requires a benign dataset) and much more efficient than FLAME (which also does not require a benign dataset). FLAME becomes very slow when there are 1000 clients in each round.
@@ 5 Related Work
Other methods find the mean or median of model update weights by excluding values based on thresholds (e.g., trimmed mean or median [33]) or frequency of occurrence (FreqFed [10]). Despite their robustness, these approaches are vulnerable to adaptive attacks, such as the Trim attack [8], which exploit these methods??? limitations.
@@ Appendix 0.B More Results on Model-boosting Attack
Model-boosting Attack on FMNIST: Figure 9 shows the results of the model-boosting attack on the FMNIST dataset. During this attack, we present the base attack with various numbers of adversaries, ranging from $1$ to $4$. It has been observed that selecting only one adversarial client per round hinders the convergence rate of the FedAvg algorithm. This is because FedAvg distributes the adversarial clients evenly based on the number of data points the model was trained on. Within a few rounds, the convergence rate drops to accuracy, and after 20 iterations, the algorithm reaches an average accuracy of around. During our experiments, we noticed a group of algorithms that gradually lost their effectiveness over time. These algorithms exhibited some fluctuations, but particularly those that ended the last 10 iterations below a target threshold of 50% caught our attention. Specifically, FLAME, Trimmed mean, FLTrust, and Krum were the affected algorithms. We suspect that the combination of non-iid ($80\%$), attack selected, and one adversary ($10\%$) among the clients could gradually decrease the performance of the aggregation algorithm over time.
Suppose we only consider the algorithms that are able to converge at a rate at or higher after 100th iterations during the one adversary per round attack. In that case, we are left with FedTruth-layer, FedTruth, FLTrust, Median, and Krum algorithms. We observed that Krum and Median consistently have a global model accuracy below the FedTruth, FedTruth-layer, and FLTrust algorithms after 100 iterations when considering the attacks that used 2, 3, or 4 adversaries per round. However, we did see Krum outperform FLTrust when there were 4 adversaries, which could be based on FLTrust selecting an inadequate representative dataset during that experiment. However, since FedTruth and FedTruth-layer use the adaptive approach, we have presented in this work that does not use a representative dataset, the results for these aggregation algorithms outperform all of the other aggregation algorithms regardless of the number of adversaries present during the experiment.
In addition, based on our observations, we noticed a similar pattern in the behavior of the FedTruth and FedTruth-layer algorithms when there were 2 to 4 adversaries per iteration. However, our results indicated that the final accuracy of the global model was better for FedTruth when there were more adversaries present (specifically, 3 or 4) compared to FedTruth-layer. This aligns with our hypothesis on the two versions of the algorithm, where the FedTruth-layer was designed to be more sensitive to minor (i.e., stealthy) model perturbations and therefore performed better when there were fewer adversaries (1 or 2).
@@
Model-boosting Attack on CIFAR-10: Figure 10 presents our results for the model-boosting attack when 3 adversaries are boosting their update by a multiple of 10 using the CIFAR-10 dataset. We significantly increased the number of iterations for the CIFAR-10 dataset experiments, conducting 1000 iterations compared to the 100 iterations used for the MNIST and FMNIST experiments. We use the ResNet-18 model when running the experiments on the CIFAR-10 datasets. After running the experiments, we noticed a slight variation in accuracy among one group of the aggregation algorithms, which include FedTruth, FedTruth-layer, FLTrust, Median, and Krum. The other group consisted of Trimmed mean and FedAvg, which exhibited poor performance with a consistent accuracy range of $0\%$ to $1\%$ throughout the training process. This was not, however, unexpected and consistent with the findings for similar attacks like Figure 10; yet for these reasons, we will not be discussing the Trimmed mean and FedAvg algorithms here.
During our experiment, we found that both FedTruth and FedTruth-layer consistently yielded the highest accuracy throughout the entire experiment. Although there was a slight increase in accuracy for our proposed algorithms when compared to FLTrust, Median, and Krum, this difference was only noticeable after the 225th iteration. Upon closer examination, after the 600th iteration, we observed a significant variance in accuracy between the Median and other algorithms. This is due to the non-iid sampling rate, causing overfitting on primary features during initial rounds; the Krum algorithm mitigates this but still affects its performance. Furthermore, we suspect that the high non-iid may also affect the FLTrust algorithm, as the weight assignment may be biased toward primary features in the server model. However, since the FedTruth and FedTruth-layer algorithms use weighted averages, new features can still be added to the model. Since the weights are computed with an adaptive truth, the influences of the chosen primary label on the clients’ devices yield minimal effects on the model’s global accuracy.

### v3

@@ 3.5 Proof of Byzantine-Resilience
BEGCuse $S$ is an arbitrary set of size $n-f$, we say that
@@ 3.6 Resisting against Adaptive Attack on FedTruth
The effectiveness of FedTruth relies on the assumption that the majority of clients are reliable and diverse. If an adversary compromises more than 50% of the clients, they can dominate the results of FedTruth. From our experimental results, we find that when an adversary compromises 40% (4 out of 10) clients in each round, FedTruth can still prevent Byzantine attacks, as seen in Figure 2. However, when the non-iid degree is further increased to 95%, as shown in Figure 7, the accuracy drops and convergence speed becomes slow bEGCuse some uncompromised clients may perform poorly with highly non-iid training data, leading to a bad estimation of the ground truth by FedTruth. However, our results outperformed all other aggregation algorithms during this experiment, excluding FLTrust.
@@ 3.7 Deploying FedTruth in Each Layer?
One major challenge in truth discovery is data heterogeneity, which may include non-structured data and missing values. However, this challenge is not applicable to FedTruth bEGCuse all the local model updates are in the same structure. For example, in deep neural networks, the model updates can be represented as multiple-layer tensors. FedTruth can be run for just one time by the aggregator to compute the truth of model updates by feeding all the local model updates into the FedTruth algorithm. This deployment treats the local model update as an observed value in the truth discovery algorithms. Also, we can deploy FedTruth on each layer to estimate the truth of that single layer, which means that the weights allocated to all the clients may vary on different layers. We denote this deployment as FedTruth-Layer. Such layer-wise deployment seems reasonable, it also brings the computation overhead which is linear to the number of layers. We compare the efficiency between FedTruth and FedTruth-Layer in the Section 4.6.
@@ 4.1 Experimental Settings
Datasets: We conduct the experiments with MNIST [7], FMNIST [30], and CIFAR-10 [16] datasets. FMNIST and MNIST are comprised of 60,000 black-and-white labeled images of size (28×28). MNIST contains handwritten digits, and FMNIST consists of images of clothing items. CIFAR-10 consists of 60000 color images of size (32x32). During the edge-case attack experiment, we used the Arkiv Digital Sweden (ARDIS) [17] dataset as the adversarial backdoor images. The ARDIS dataset consists of handwritten digits originating from Swedish church records. This dataset is suitable for targeted images when inserting backdoors into MNIST, as ARDIS entirely consists of naturally occurring edge cases.
Clients: When crafting the clients’ local datasets, we draw their datapoints randomly from a non-iid distribution. We use a non-iid distribution bEGCuse it better represents clients’ data in practice than an iid distribution. The client’s local data is generated a non-iid distribution, where the bias parameter default is $0.8$. In addition, we evaluate the impact of non-iid degree using the model-boosting attacks, as seen in Section 4.4. In each FL round, we randomly select 10 clients and choose a subset of these selected clients as adversarial clients.
@@ 4.2 Byzantine Attacks
Figure 2(a) shows how all of the aggregation algorithms perform when there are no adversaries present. Figure 2(b) presents the results when $10\%$ of the clients are compromised in each round. We observe that all of the aggregation algorithms performed well except for FedAvg. However, when the percentage of compromised clients increases to 30%, as shown in Figure 2(c), the FedTruth methods remain unaffected by the attack. However, Trimmed-mean, similar to FedAvg, is significantly impacted at this stage and does not reach convergence.
The results in Figure 2(d) show that increasing the number of adversaries per iteration to 40% slows the convergence rate for the FedTruth, FedTruth-Layer, and Median aggregation algorithms. However, they are still able to reach an accuracy of 80% after the 100th iteration. The FedAvg and Trimmed-mean algorithms were compromised during this version of the experiment as well, preventing them both from reaching any convergence when at least $20\%$ of the clients are adversarial. In contrast, the remaining algorithms (FLTrust, Krum, FLAME) were not affected during this attack, regardless of the number of adversaries we selected.
Gaussian Noise Attack: The Gaussian noise attack aims to degrade the performance of the global model by adding arbitrary noise to the model. The noise is drawn from a multivariate Gaussian distribution $N(0,\sigma^{2}I)$ [5, 4] and is added directly to the adversaries’ model before sending it to the aggregator.
@@
In Figure 3(a), we show the accuracy and convergence speed of the model for all the aggregation algorithms against the Gaussian noise attack, where three adversaries launch this attack per round. Our findings are as follows: 1) FedTruth and FedTruth-Layer can defend against the Gaussian noise attack without significantly slowing down the convergence speed; 2) FedTruth and FedTruth-Layer can achieve the same model accuracy as FLTrust, which requires a benign dataset, showing that our proposed algorithm can actually find the ground truth of the model updates; and 3) FedAvg cannot converge within 100 rounds under the Gaussian noise attack.
@@ 4.3 Backdoor Attacks
In this section, we present our findings for target task accuracy (accuracy on an adversarial backdoor dataset) and the main task accuracy (accuracy on a benign dataset) for the distributed backdoor attack (DBA) during the base,model-boosting, and constrain-and-scale attacks. During these attacks, we used the cosine distance metric, as it presented the best results during the majority of backdoor attack configurations. We provide results for the projected gradient descent and edge-case attacks with the same configurations in Appendix 0.D and analysis on the effects of different distance metrics in Section 4.5 and Appendix 0.E.
@@
Figures 4(a), 5(a), and 6(a) show the main task accuracy for all of the versions of the DBA attack. Here, we see that both Krum and Flame are unable to train the main task during the base attack and when combined with the model-boosting. We suspect these results could be caused by the large number of adversaries colluding during this experiment or the imbalanced sample data. The remaining algorithms are able to reach convergence on the main task during these attacks. When the DBA is combined with the constrain-and-scale attack, all algorithms are able to reach convergence, including Flame and Krum, which we suspect is due to the constrain-and-scale attack reducing the amount that an adversarial model can diverge during each epoch. However, FedTruth and FedTruth-Layer experience a slower convergence rate during the DBA with model-boosting and constrain-and-scale attacks.
@@
Figures 4(b), 5(b), and 6(b) present our findings on the targeted task accuracy (i.e., the backdoor accuracy) during the DBA experiments. The adversarial goal during this attack is to add the targeted artifact to the global model without being detected and without affecting the model’s performance on the main task. Therefore, we will not be considering the Krum and FLAME algorithms during the base and model-boosting attacks, as they were unable to converge when training the main task, as seen in Figure 4(a) and 5(b).
During the base attack (Figure 4(b)), FedTruth and FedTruth-Layer are the only aggregation algorithms that are able to remove the backdoor, finishing with a backdoor accuracy below 5%. Furthermore, during the DBA with model-boosting attack, we observe similar results with FedTruth and FedTruth-Layer being able to remove the backdoor and reach a final backdoor accuracy below 5%. However, FLTrust is now able to reach a similar convergence rate, and the Median algorithm improves as well, with a final accuracy below 40%. This is a result of a tradeoff between the stealthiness of the base DBA attack being diminished when increasing the amplitude of the adversarial updates vector in hopes of replacing the global model with the adversarial model. Our results also show that the DBA with constrain-and-scale attack is effective at adding the targeted task into FedAvg, FLTrust, Median, and Trimmed mean, reaching a backdoor accuracy above 40% after 100 iterations. However, FedTruth, FedTruth-Layer, Krum, and FLAME are able to remove the adversarial artifact finishing with a backdoor accuracy below 5%.
@@ 4.4 The Impact of non-iid on FedTruth
Figure 7 shows how various non-iid bias parameters affect the experiments when adversaries apply the model-boosting attack. During these experiments, three adversaries were selected in each round. Our methodology for sampling non-iid data was specifically engineered to replicate varying degrees of label bias, thus enabling an in-depth analysis of its influence on federated learning model efficacy. We manipulated a bias parameter to adjust the label proportions within each client’s local dataset. For instance, setting the bias parameter to 0.9 indicated that 90% of a client’s dataset contained instances of their primary label, with the remaining 10% consisting of instances from other labels, allocated based on a Gaussian distribution. For this experiment, we set the bias parameters as $0.1$, $0.5$, $0.8$, and $0.95$.
@@ 4.5 Distance Function in FedTruth
From the FedTruth formulation (i.e., Equ. 1), we can see that the distance function plays a significant role in separating benign and malicious model updates. This section discusses how FedTruth performs with different distance functions. More results on different distance metrics of FedTruth will be shown in the Appendix 0.E.
@@
Figures 8(a) and 8(b) show how the choice of distance metric affects FedTruth during Byzantine attacks. As expected, the two metrics that solely measure the difference between two points (Euclidean and Manhattan distances) performed the best, reaching convergence in both cases with a final accuracy of 100%. This is because these approaches can easily identify the adversarial model furthest from the benign models, as these are either boosted or have additional random noise inserted into them. During the Model Boosting attack (Figure 8(a)), we see that the custom distance metric was able to reach sub-optimal performance, finishing with an accuracy of 60%. However, since the angular distance and cosine distance metrics do not consider a model’s magnitude, they are ineffective during the model boosting attack (Figure 8(a)). During the Gaussian noise attack (Figure 8(b)), the angular, cosine, and custom distances were also not effective. This is due to the slight modification of the angle during the Gaussian noise attack, causing the adversarial updates to be less effective. This is also why the custom distance, which partially relies on the Euclidean distance, is not sufficient for removing the adversarial updates.
Figures 8(c) and 8(d) present our results for all distance metrics during a backdoor (DBA) attack. These results indicate that the cosine distance metrics perform the best, as it is able to remove any backdoors injected into the model. Interestingly, while the angular distance finished with a final backdoor accuracy of 60%, while the cosine distance metric was able to reduce the backdoor accuracy to below 5%. The models focused on point-to-point distance measurement (Euclidean, Manhattan, and custom distance) did not perform well, with all of their final backdoor accuracies reaching 100%.
@@ 4.6 Efficiency Evaluation of FedTruth and FedTruth-Layer
We find that during the MNIST experiments, the Median algorithm is the most efficient, while Trimmed mean, FedAvg, and FLTrust take less than 0.02 seconds. FLAME takes about 0.16 seconds to filter and clip, while Krum takes around 0.08 seconds. Interestingly, FedTruth has a slightly faster aggregation time than FLAME, and FedTruth-Layer is the slowest aggregation algorithm with an average aggregation time of 0.7 seconds.
@@
We count the number of iterations required for both algorithms to reach convergence and present the results in Table 2. For the CNN (8 layers) model on the MNIST dataset, we find that FedTruth requires an average of 5.17 iterations to estimate the ground-truth model update, while FedTruth-Layer requires an average of 33.73 iterations (six times more than FedTruth) to reach convergence, despite having eight layers in the CNN model. Moreover, FedTruth on each layer has a smaller input size and requires less computation time on the distance compared to FedTruth with the entire model update as input. In conclusion, we find that both FedTruth and FedTruth-Layer have similar performance on the MNIST CNN (8 layers) model regarding accuracy and robustness. However, FedTruth is much more efficient than FedTruth-Layer when the number of layers increases, which can be observed in the CIFAR-10 (ResNet-18) model.
We further evaluate the average aggregation time for 10, 100, and 1000 clients in a single FL round in Table 3, where the aggregation time is calculated as the average of 100 FL rounds. We can see that FedTruth and FedTruth-Layer are as efficient as FLTrust (which requires a benign dataset) and much more efficient than FLAME (which does not require a benign dataset). FLAME becomes very slow when there are 1000 clients in each round.
@@ 5 Related Work
Other methods find the mean or median of model update weights by excluding values based on thresholds (e.g., trimmed mean or median [33]) or frequency of occurrence (FreqFed [10]). Despite their robustness, these approaches are vulnerable to adaptive attacks, such as the Trim attack [8], which exploit the limitations of these methods.
@@ Appendix 0.B More Results on Model-boosting Attack
Model-boosting Attack on FMNIST: Figure 9 presents the results of the model-boosting attack on the FMNIST dataset, evaluating the impact of varying the number of adversaries per round from $1$ to $4$.
Figure 9(a) illustrates that when one adversarial client is selected per round, FedAvg is prevented from reaching convergence. This is likely due to FedAvg’s proportional distribution of adversarial updates, which, under non-iid data conditions, can amplify the effect of even a single adversary. As a result, FedAvg’s accuracy rapidly declines, stabilizing at an accuracy of 0%. A subset of algorithms, specifically FedTruth-Layer, FLAME, Trimmed Mean, FLTrust, and Krum, also exhibit noteworthy degradation, with their performance reaching a final accuracy below 50%. This trend suggests that the combination of non-iid data (80%), the attack configuration, and the presence of even a single adversary (10% of clients) can undermine the robustness of these aggregation strategies over time. However, FedTruth and Median offer the best results, reaching a convergence rate above 75%.
Figures 9(b)-9(d) show that with 2, 3, or 4 adversaries, neither Median nor FedAvg converges. Furthermore, Krum reaches a convergence rate of 35% when 3 adversaries are present, as seen in Figure 9(c). The remaining algorithms during these experiments reach a minimum convergence rate of above 40%. Interestingly, when 4 adversaries are present (Figure 9(d)), FLTrust is the worst performing algorithm that reached convergence. This is potentially due to FLTrust’s reliance on a representative root dataset, which may have been suboptimal. In contrast, FedTruth and FedTruth-Layer, which do not depend on a representative dataset and instead use adaptive aggregation, consistently outperform all other methods regardless of the number of adversaries.
We observe that FedTruth and FedTruth-Layer display similar robustness patterns as the number of adversaries increases with one or both reaches an optimal convergence rate. However, FedTruth achieves higher final accuracy than FedTruth-Layer when more adversaries are present (specifically, 2 - 4). This aligns with our hypothesis that FedTruth-Layer, designed for heightened sensitivity to subtle (stealthy) attacks, excels with fewer adversaries and, while FedTruth is more resilient as attack intensity increases.
@@
Model-boosting Attack on CIFAR-10: Figure 10 presents our results for the model-boosting attack, in which between $1$ and $4$ adversaries per iteration boost their local updates by a factor of $10$ on the CIFAR-10 dataset. We increased the number of iterations to 1,000, compared to the 100 iterations used for the MNIST and FMNIST experiments, to accommodate the slower convergence rate observed under our hyperparameter settings and non-iid sampling constraints. We used the ResNet-18 model for all CIFAR-10 experiments.
Figure 10(a) presents results for the model-boosting attack on the CIFAR-10 dataset when 1 adversary is present. These results indicate that all algorithms except FedAvg and FLAME converge, with both algorithms’ accuracy falling below 1% after 1,000 iterations. Krum’s accuracy reaches approximately 40%, lower than the other algorithms which reached convergence. Trimmed mean, FedTruth, and FedTruth-Layer’s accuracy exceed 60%, while the remaining algorithms finish above 50%.
Figures 10(b) and 10(c) show our results with 2 and 3 adversaries, respectively, revealing similar trends to Figure 10(a). Notably Trimmed mean fails to converge in both cases. When there are 2 adversaries, the Median algorithm’s accuracy falls to just below 40%, and with 3 adversaries, it falls further to around 30%. In contrast, FedTruth and FedTruth-Layer maintain high final accuracy, unaffected by the increased number of adversaries. FLTrust converges to about 50% accuracy after 1,000 iterations in the presence of 3 adversaries.
Figure 10(d) reports our results when there are 4 adversaries during each aggregation round. In this experiment, the Median algorithm’s accuracy collapsed, plateauing at 1%, significantly below our proposed algorithms’ 60% final accuracy. FLTrust’s accuracy drops below Krum’s, with both ending slightly below 40%. In contrast, FedTruth and FedTruth-Layer maintain a robust convergence rate.
Overall, FedTruth and FedTruth-Layer consistently achieve final accuracies above 60% across all attack configurations. As adversaries increase, their performance advantage over FLTrust, Median, and Krum becomes more pronounced, with divergences observable after iteration 250 in Figure 10(b), 225 in Figure 10(c), and 200 in Figure 10(d). The diminished performance of Trimmed Mean, Krum, and Median can be attributed to their higher likelihood of incorporating adversarial or non-representative updates, which destabilizes aggregation. Krum, in particular, may overfit to dominant features early on due to non-iid sampling, yielding a suboptimal global model. FLTrust appears susceptible to bias toward server-side features when benign client contributions are limited. In contrast, FedTruth and FedTruth-Layer employ adaptive weighted averaging, enabling the integration of new features and minimizing the influence of any single label, regardless of the number of adversaries present.